### v1

# PLaID++: A Preference Aligned Language Model for Targeted Inorganic Materials Design

Abstract
Discovering novel materials is critical for technological advancements such as solar cells, batteries, and carbon capture. However, the development of new materials is constrained by a slow and expensive trial-and-error process. To accelerate this pipeline, we introduce PLaID++, a Large Language Model (LLM) fine-tuned for stable and property-guided crystal generation. We fine-tune Qwen-2.5 7B to generate crystal structures using a novel Wyckoff-based text representation. We show that generation can be effectively guided with a reinforcement learning technique based on Direct Preference Optimization (DPO), with sampled structures categorized by their stability, novelty, and space group. By encoding symmetry constraints directly into text and guiding model outputs towards desirable chemical space, PLaID++ generates structures that are thermodynamically stable, unique, and novel at a $\sim$50% greater rate than prior methods and conditionally generates structures with desired space group properties. Our experiments highlight the effectiveness of iterative DPO, achieving $\sim$115% and $\sim$50% improvements in unconditional and space group conditioned generation, respectively, compared to fine-tuning alone. Our work demonstrates the potential of adapting post-training techniques from natural language processing to materials design, paving the way for targeted and efficient discovery of novel materials.
1 Introduction
The discovery of new solid-state materials is the foundation of many transformative technologies including solar cells (Green et al., 2014), batteries (Zhao et al., 2020), and carbon capture (Sriram et al., 2024a). However, the search for new materials is constrained by the immense scale of chemical space—previous explorations have only uncovered a fraction of the total number of potential stable inorganic compounds (Davies et al., 2016). Generative models offer a promising avenue for accelerating technological breakthroughs by efficiently discovering novel and stable structures in unexplored regions of chemical space.
Previous work has applied variational autoencoders (Xie et al., 2021) and denoising models (Zeni et al., 2025; Jiao et al., 2023; Miller et al., 2024) to generate stable and novel structures. However, these works either do not explicitly encode crystallographic symmetry or use computationally inefficient and complex representations, limiting the efficacy of their models.
Meanwhile, language models (Gruver et al., 2024; Sriram et al., 2024b) have emerged as a promising alternative. Their pretrained knowledge makes them data-efficient, often requiring far fewer domain-specific examples than training models from scratch (Gruver et al., 2024). Moreover, their natural-language interface allows one model to be applied to many tasks, including unconditional and conditional generation, infilling, and crystal structure prediction (Gruver et al., 2024).
Symmetry is a defining aspect of crystal structures. The set of rotations, reflections, inversions, and translations exhibited by a crystal lattice form its space group. A crystal’s space group and symmetries are not merely a mathematical construct but critical to many optical, electrical, and magnetic properties like piezoelectricity (Malgrange et al., 2014; Yang et al., 2005). One way to define a crystal’s structure from its symmetries is via Wyckoff positions, whereby one can specify a few key atomic coordinates and have the remaining atomic positions filled by applying symmetry operations (Hahn et al., 1983).
Other fields of machine learning have shown that general-purpose architectures that focus on scale and data expressivity outperform models with handcrafted, domain-specific constraints (Dosovitskiy, 2020; Kaplan et al., 2020). We contend that the same “Bitter Lesson” holds true for materials generation (Sutton, 2019). Rather than imposing explicit crystallographic constraints, we train an LLM to learn and exploit structural parameters from the data itself. By allowing the model to discover patterns implicitly through training on Wyckoff-based text representations, we see significant improvements in generation performance.
To guide our search space towards chemically useful structures, we introduce Reinforcement Learning from Interatomic Potentials (RLIP), a reinforcement learning framework for physically grounded materials generation. To do so, we adapt Direct Preference Optimization (DPO) (Rafailov et al., 2023), a computationally efficient reinforcement learning method. Our pipeline is as follows: first, we perform supervised finetuning on a base LLM with Wyckoff-based text encodings of crystals. Then we further fine-tune the LLM via multiple rounds of DPO on generated structures categorized by their stability, novelty, and space group. To prevent entropy collapse and increase the diversity of generated structures, we increase the sampling temperature across successive iterations of DPO.
By incorporating inherent crystal symmetries and the feedback of a Machine Learning Interatomic Potential (MLIP), we significantly increase the rate of stable, unique, and novel (S.U.N.) materials. Our experiments demonstrate that our method PLaID++ generates materials at state of the art stability rate and generates S.U.N materials at over a 20% higher rate than prior models while retaining the flexibility of natural language prompting. We also highlight the effectiveness of reinforcement learning for materials discovery, achieving $\sim$115% and $\sim$50% improvements in unconditional and space group conditioned generation, respectively, compared to fine-tuning alone. Our contributions are as follows:
  1. 1.
    We develop a novel, symmetry-informed text representation for crystal structures which is compact, performant, and physically-motivated.
  2. 2.
    We introduce an end-to-end framework for fine-tuning LLMs for material generation using Reinforcement-Learning from Interatomic Potentials (RLIP).
  3. 3.
    We demonstrate that fine-tuned LLMs outperform existing state-of-the-art models in generating novel and stable materials.
Figure 1: Overview of the PLaID++ pipeline, highlighting Wyckoff fine tuning and iterative DPO.
2 Related Work
Variational autoencoders (VAEs) were among the early approaches to crystal structure generation, using learned latent spaces to produce physically plausible atomic arrangements (Xie et al., 2021; Court et al., 2020). Specifically, CDVAE introduced a continuous representation of crystal structures, however, VAEs face challenges in generating structures with strict crystallographic constraints, as their representations often lack direct symmetry considerations.
Diffusion models like DiffCSP and FlowMM iteratively denoise atomic positions to generate crystal structures, achieving higher stability and validity than earlier VAE approaches (Jiao et al., 2023; Miller et al., 2024). While these models excel at capturing complex chemical landscapes, they face significant inference cost due to multi-step denoising, which limits their scalability for rapidly screening a broad sample of material candidates.
Language models are scalable and capable when leveraging extensive pre-training on vast text corpora. In particular, CrystalLLM (Gruver et al., 2024) pioneered the use of fine-tuned LLMs for crystal generation, showcasing the effectiveness of text-based representations for materials prediction.
Symmetry-aware language models have introduced Wyckoff-position-based representations, leveraging inherent crystallographic symmetries to more effectively constrain the generative process within physical reality. WyFormer and CrystFormer (Kazeev et al., 2024; Cao et al., 2024) utilized transformer-based architectures to predict atomic positions conditioned explicitly on predefined space groups and chemical formulas. Although these symmetry-aware methods showed improvements in structural validity, their reliance on explicit formula and space group constraints considerably limits their utility in exploratory de novo generation tasks.
On the other hand, symmetry-preserving non-LLM methods such as WyCryst (Zhu et al., 2024), Crystal-GFN (AI4Science et al., 2023), DiffCSP++ Jiao et al. (2023), and SymmCD Levy et al. (2024) enforce Wyckoff or space group constraints through templates or diffusion processes. While effective, these approaches either require predefined templates or incur high computational cost, and cannot flexibly target multiple objectives. In this paper, we explore using a compact Wyckoff text encoding with preference-based post-training, enabling both unconditional and conditional generation in a single unified pipeline.
3 Background
3.1 Crystallography
Crystal structures are periodic arrangements of atoms defined by repeating unit cells. Each unit cell is fully described by lattice parameters—consisting of side lengths $(l_{1},l_{2},l_{3})$ and angles $(\theta_{1},\theta_{2},\theta_{3})$—and atomic positions within the cell, characterized by their fractional coordinates $(x_{i},y_{i},z_{i})$ and elemental identities $e_{i}$. Formally, a crystal structure $C$ can be expressed as a tuple:
$C=(l_{1},l_{2},l_{3},\theta_{1},\theta_{2},\theta_{3},e_{1},x_{1},y_{1},z_{1},\dots,e_{N},x_{N},y_{N},z_{N}).$ (1)
The properties of crystals are heavily influenced by their underlying symmetry, captured mathematically by their space groups. Each space group comprises a set of symmetry operations, such as rotations, reflections, and inversions, that map the crystal structure onto itself. Atomic positions consistent with these symmetry operations are defined through Wyckoff positions (Hahn et al., 1983), significantly reducing the degrees of freedom required to specify a structure. Consequently, employing Wyckoff positions in generative modeling can greatly enhance the validity and physical plausibility of generated structures. We present additional mathematical background on Wyckoff positions in Appendix 7.5.
In computational materials science, thermodynamic stability is a critical metric which can be described using the energy above hull ($E^{\text{hull}}$), which quantifies how energetically favorable a material is compared to known competing phases (Sun et al., 2016). The energy hull refers to a convex hull of formation energies in composition-energy space constructed by evaluating the formation energy of each computed phase and linking the lowest-energy points to form the lower convex envelope (Gruver et al., 2024). Hence, the $E^{\text{hull}}$ of a candidate material is defined as the distance between its computed formation energy per atom and the hull at the same composition.
Materials with $E^{\text{hull}}\leq 0$ eV/atom at a given temperature and pressure are considered thermodynamically stable, while those slightly above zero (typically $E^{\text{hull}}<0.1$ eV/atom) are considered metastable and potentially synthesizable (Jain et al., 2013). Compounds with lower $E^{\text{hull}}$ values have less thermodynamic driving force for decomposition, making them generally more likely to be synthesizable and thus have potential for real-world applications. Consequently, many generative models prioritize designing materials with low $E^{\mathrm{hull}}$ to maximize their physical viability.
Crystallography concerns not only the validity of atomic positions and crystal stability but also the uniqueness and novelty of generated structures. In practice, two crystals may have different symmetries yet could be considered chemically equivalent if their lattices or atomic sites overlap under tolerance thresholds. To address this, we rely on structure-based fingerprints such as Pymatgen’s StructureMatcher (Ong et al., 2013). StructureMatcher reduces candidate crystals to canonical forms (e.g., primitive and Niggli cells) and then checks atom-by-atom equivalence within length, angle, and site tolerances. Uniqueness and novelty therefore depend on the chosen reference set and tolerance hyperparameters. These metrics underpin the S.U.N. evaluation used throughout this work, complementing stability as a necessary criterion for materials discovery.
3.2 Reinforcement learning for LLM fine-tuning.
Recent advances in reinforcement learning have enabled LLMs to align generated outputs more closely with human preferences (Bai et al., 2022). In general, these methods fine-tune a pre-trained LLM, $\pi_{ref}$ with feedback from a reward model $r_{\phi}(\cdot)$. The RL objective optimizes the model $\pi_{\theta}(y\mid x)$ to maximize the expected reward of generated responses $y\sim\pi_{\theta}(y\mid x)$, given prompt $x$. To preserve the diversity of response and prevent collapse to undesirable modes of the reward function, this is often regularized via an additional KL-divergence loss that prevents the model from straying too far from the initial model. This leads to the following general objective:
$\max_{\pi_{\theta}}\mathbb{E}_{x\sim D,y\sim\pi_{\theta}(y\mid x)}\left[r_{\phi}(x,y)\right]-\beta D_{\mathrm{KL}}\left[\pi_{\theta}(y\mid x)\|\pi_{\text{ref}}(y\mid x)\right].$ (2)
Here $D$ represents the distribution of prompts and the hyperparameter $\beta$ controls the strength of regularization. This objective can be optimized using REINFORCE or with reduced variance policy gradient such as proximal policy optimization (PPO) Schulman et al. (2017). This framework requires an explicit reward function, while actor-critic methods such as PPO additionally require a value model leading to substantially increased computational overhead.
In contrast, Direct Preference Optimization (DPO) (Rafailov et al., 2023) streamlines this process, using observed preference pairs of responses $(y_{w},y_{l})$, generated from $\pi_{ref}$, where $y_{w}$ denotes the preferred response. DPO optimizes an implicit reward function $r^{*}(x,y)$ by maximizing the likelihood of preferences under a Bradley-Terry preference model: $p(y_{w}\succ y_{l}\mid x)=\exp(r^{*}(y_{w}|x))/(\exp(r^{*}(y_{w}|x))+\exp(r^{*}(y_{l}|x)))$ . The authors show that a KL-regularized objective can be optimized without directly instantiating $r^{*}(x,y)$, leading to the following objective:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(x,y_{w},y_{l})}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\thinspace|\thinspace x)}{\pi_{\text{ref}}(y_{w}\thinspace|\thinspace x)}-\beta\log\frac{\pi_{\theta}(y_{l}\thinspace|\thinspace x)}{\pi_{\text{ref}}(y_{l}\thinspace|\thinspace x)}\right)\right]$ (3)
Prior work has shown that DPO is effective in both RLHF (Rafailov et al., 2023) and scientific domains (Widatalla et al., 2024).
To effectively generate stable structures, we employ a reinforcement learning framework based on DPO. By curating a dataset of stability and novelty-ranked crystal pairs evaluated via MLIP relaxations, our model iteratively learns to favor generation of structures with lower predicted energies in unseen crystal space, thus aligning generative behavior towards chemically stable and experimentally realizable new materials.
Figure 2: Left: An example crystal highlighting the $p4mm$ space group symmetry, where colors represent atoms of different elements. We represent the entire asymmetric unit of the crystal in our Wyckoff-based text representation by leveraging the crystal’s symmetry. Right: The template from which we generate prompts used for training. For conditional generation, we include the blue conditioning information, and for unconditional generation, we remove it from our prompt. In all prompts, the crystal string is replaced with the encoding on the left.
4 Method
4.1 Base Model
Our approach begins with a pre-trained large language model (Qwen-2.5 7B) (Yang et al., 2024), which has demonstrated strong capabilities in text generation and reasoning. Following the methodology outlined by (Gruver et al., 2024), we fine-tune the base version of Qwen-2.5 7B to model the crystal generation process via parameter-efficient fine-tuning (Hu et al., 2021). Unlike prior works that rely on explicit space group constraints or predefined templates, our method enables more flexible and generalizable unconditional and conditional structure generation through the use of a natural language prompt.
The training objective is to maximize the likelihood that the model generates the correct token sequence representing stable crystal structures. We fine-tune two versions of the base model: one where the crystal structures are represented as text-based 3D coordinates and another based on a new text-based Wyckoff representation.
4.2 Wyckoff Representation
Crystal structures inherently follow symmetry constraints that can be represented by space groups and Wyckoff positions. To improve the model’s ability to generate valid and stable structures, we introduce a Wyckoff representation that encodes these symmetries directly into text. Unlike previous methods such as DiffCSP++ (Jiao et al., 2024) which rely on searching template structures for Wyckoff positions, our approach allows direct prediction of these sites within the generative process.
Because Wyckoff positions specify atomic site symmetries in a given space group, they significantly increase the possibility of feasible atomic arrangements. Our encoding method represents each crystal as a sequence of tokens as shown in Figure 2. The prompt used to train PLaID++ is the same for both unconditional and conditional generation except conditional generation includes the conditioning information on space group in blue. We specify the chemical formula, space group, lattice parameters, individual chemical elements, fractional coordinates, and Wyckoff site symmetries. Note that we include the chemical formula first to ensure self-consistency during generation.
4.3 Direct Preference Optimization (DPO)
Encoding symmetry improves structural validity; however, stability, uniqueness, and novelty remain critical for practical material discovery. To explicitly align our model toward generating crystals with desired properties, we employ an iterative form of Direct Preference Optimization (DPO) (Rafailov et al., 2023), a reinforcement learning technique that directly optimizes model outputs based on ranked preferences as highlighted by Equation 3. We jointly optimize our model on three explicit objectives: thermodynamic stability, novelty with respect to a reference crystal set, and space group conditioning.
Stability   To train the model to prefer stable crystals, we use EquiformerV2 86M (eqV2) (Barroso-Luque et al., 2024) as a reward signal. eqV2 is a Machine Learning Interatomic Potential (MLIP) that predicts relaxed formation energies for materials, serving as an efficient proxy for stability assessment. To guide the model towards generating more stable materials, we categorize a crystal’s energy above the hull into one of three buckets: stable ($\leq 0\text{ eV/atom }E^{\text{hull}}$), metastable ($\leq 0.08\text{ eV/atom }E^{\text{hull}}$), or unstable ($0.08\text{ eV/atom }E^{\text{hull}}$). We create stability preference pairs $(x,y_{w},y_{l})$, where $x$ is the model prompt and $y_{w}$ and $y_{l}$ are accepted and rejected crystals. Concretely, we sample preference pairs to form (stable, metastable) and (stable/metastable, unstable) sets, forming a tiered dataset that encodes relative preferences over increasing degrees of thermodynamic stability. This allows us to incorporate fine-grained (stable, metastable) reward signals into our training set without overfitting to exact eqV2-predicted $E^{\text{hull}}$ values and without relying heavily on stable generations from the base model, which are too infrequent to provide sufficient training signal as seen in Appendix 7.2.
Uniqueness and Novelty   To incorporate novelty into our reward formulation, we construct preference pairs that distinguish between (stable & novel, stable & not novel). Incorporating uniqueness into a pairwise reward is non-trivial and presents many challenges. Namely, since uniqueness is defined at the group level (i.e. whether or not crystals within a set of rollouts are considered the “same” crystal as defined by Pymatgen’s StructureMatcher), pairwise reward computation of uniqueness is ambiguous. However, we can control sample diversity during rollouts via the sampling temperature. As temperature increases, uniqueness and novelty tend to increase, while stability and validity of generated crystals decrease. Based on this observation, we find that increasing the sampling temperature across successive iterations of DPO can prevent overoptimization for stability, preventing the model from collapsing on narrow regions of chemical space.
Space Group   To condition our model on generating crystals with a desired space group, we add an additional constraint during preference pair construction. For each space group, we use a conditional generation prompt indicating the target space group number and compare crystals that match this group. Preference pairs are then formed by contrasting (1) stable or metastable crystals with the correct space group against other stable or metastable crystals with the incorrect space group, and (2) stable or metastable structures against unstable structures. This strategy requires stability as a base condition for generation, but also encourages the model to align for structural symmetry.
Fine-Tuning   We emphasize that both the S.U.N. and space group preference pairs are jointly used in the same DPO objective across the unconditional prompt and space group specific conditional prompts. This joint optimization encourages the model to learn to generate both generally stable crystals and space-group-specific structures in a unified pipeline.
To generate this dataset, we sample 10,000 crystal structures from our fine-tuned Qwen model for unconditional generation and sample 1,000 crystal structures across each of seven space groups. More details about our space group conditional generation is provided in Section 5.1; From here, we categorize our samples into our preference datasets based off the above metrics, resulting in a universal dataset that covers both unconditional and conditional generation.
We apply DPO to fine-tune Qwen on this curated preference dataset. We adopt an iterative DPO approach in which $\pi_{\text{ref}}=\pi_{\theta-1}$, with generations from $\pi_{\theta-1}$ used to construct the set of preference pairs for $\pi_{\theta}$. We have additional information about our DPO hyperparameters, our dataset creation process, as well as other potential DPO variants and ablations in Appendix 7.2.
5 Experiments
5.1 Setup
We trained our model on the well-established MP-20 dataset (Xie et al., 2021), a collection of 45,231 inorganic crystalline materials from the Materials Project (Jain et al., 2013). The dataset includes structures with up to 20 atoms, all of which are metastable. We follow the methodology of Gruver et al. (2024) by independently fine-tuning a pre-trained Qwen-2.5 7B model using 4-bit quantization and Low-Rank Adapters (LoRA) (Hu et al., 2021), implemented with PyTorch (Paszke et al., 2019) and Transformers (Wolf et al., 2020). Symmetry information for our Wyckoff representation is calculated by Pyxtal Fredericks et al. (2021). Following supervised fine-tuning, we apply DPO on the generated preference dataset to further guide generation towards stable structures. Full hyperparameters details are provided in Appendix 7.2.
For unconditional generation, we sample 10,000 structures from each fine-tuned model, parsing a CIF from the generated string. We resample if a CIF cannot be parsed from the string, which guarantees all samples can be interpreted as crystals but does not guarantee validity.
Similarly, for space-group conditioned generation, we sample 1,000 structures for each of seven space groups. These space groups were chosen in accordance with (Zeni et al., 2025) as they represent each of the seven crystal systems with varying levels of symmetry. Specifically, we sample from space groups $P1,\ C2/c,\ Amm2,\ I\overline{4}m2,\ P3,\ P6_{3}/mmc,\ \mathrm{and}\ F\overline{4}3m$.
To validate the stability and accuracy of our MLIP-based results, we performed Density Functional Theory (DFT) calculations on crystal samples generated by our flagship Qwen-based model fine-tuned with DPO and the Wyckoff encoding. Although accurate, DFT is computationally expensive, scaling typically as $O(N^{3})$ with respect to the number of atoms $N$. Therefore, in this work, in addition to performing DFT calculations, we also leverage machine learning interatomic potentials (MLIPs) such as eqV2 (Barroso-Luque et al., 2024) and eSEN (Fu et al., 2025) as efficient proxies for DFT. These MLIPs predict the relaxed atomic positions and energies with significantly reduced computational overhead. We specifically use eSEN for evaluation to ensure our generative models are guided by energetics closely aligned with DFT benchmarks, without incurring DFT’s computational expense during training. To verify our eSEN results, we additionally take a random 1,000 crystal subset of our generated crystals to compute stability using DFT. Detailed information regarding our DFT setup are in Appendix 7.2.
Table 1: Results for unconditional materials generation on the MP-20 dataset. Our flagship PLaID++ variant uses a Wyckoff text representation, DPO on stability and novelty-based preference pairs, and dynamic temperature adjustment. Column-best figures are in bold.
Method Params Validity (%) (↑) Stability (%) (↑) S.U.N. (%) (↑)
Structural Composition
CDVAE 100.00 86.70 1.6
DiffCSP 100.00 83.25 5.0 3.3
FlowMM 96.85 83.19 4.6 2.8
FlowLLM 99.94 90.84 13.9 4.7
MatterGen-MP 13.0
CrystalLLM-70B 99.60 95.40 5.28
Jointly-trained ADiT 99.74 92.14 15.4 5.3
PLaID++ (Non-DPO Types) 3D Coord 98.93 88.94 7.20 2.81
Wyckoff 99.80 91.22 7.17 3.58
Iterative SFT 99.85 94.54 9.85 4.13
PLaID++ (DPO Types) 3D Coord 96.62 89.94 13.94 1.69
Wyckoff 98.81 92.57 15.59 6.25
Stability + Novelty 99.78 95.64 16.72 6.22
PLaID++ 99.75 97.34 22.27 7.74
5.2 Metrics
To initially evaluate the quality of our generated crystal structures, we focus on structural and compositional validity, as defined by Xie et al. (2021). These metrics provide an effective proxy for assessing the quality of generated crystals before conducting more computationally expensive stability evaluations. We explain more details about these metrics in Appendix 7.1
Our primary metric for evaluating generated crystal structures is the S.U.N Rate from Miller et al. (2024), which measures the percentage of crystals that are stable, unique, and novel. Stability is assessed by comparing a crystal’s energy to a convex hull of previously computed energies from Riebesell et al. (2023). Crystals on or below the hull ($\leq 0\text{ eV/atom }E^{\text{hull}}$) are deemed stable. We evaluate uniqueness—differentiation from other generated crystals—and novelty, which measures diversity from the training data to calculate S.U.N.
For our space group conditional generation task, we adopt the S.S.U.N Rate metric from Zeni et al. (2025), which measures the percentage of strucutures that are the correct symmetry group as deemed by Pyxtal (Fredericks et al., 2021) and are metastable ($\leq 0.1\text{ eV/atom }E^{\text{hull}}$), unique and novel.
Due to compute constraints, we assess stability for unconditional and conditional generation primarily via the eSEN MLIP (Fu et al., 2025). Note that we specifically use different MLIPs—eqV2 and eSEN—for preference dataset creation and crystal generation evaluation to avoid reward hacking or overfitting to a specific MLIP.
Figure 3: Evolution of S.U.N. percentage of PLaID++ variants’ over DPO iterations. Reference lines represent S.U.N. rates from ADiT and FlowLLM’s flagship models.
5.3 Results
Unconditional Generation   We compare our model to seven prior methods, as reported from Joshi et al. (2025): CDVAE (Xie et al., 2021), DiffCSP (Jiao et al., 2023), FlowMM (Miller et al., 2024), FlowLLM (Sriram et al., 2024b), MatterGen-MP (Zeni et al., 2025), CrystalLLM (Gruver et al., 2024), and ADiT (Joshi et al., 2025). Our main results are presented in Table 1. On stability and S.U.N, the most important metrics, PLaID++ outperforms all prior methods. We also analyze the contribution of each training technique to our methodology. We see that our novel Wyckoff representation and iterative DPO both significantly improve S.U.N., resulting in our best PLaID++ model. For our best overall PLaID++ model, 22.27% of the relaxed structures are deemed stable, out of which 72% are novel, and 55% are unique, leading to a S.U.N. rate of 7.74%. PLaID++ achieves the highest stability rate and a $\sim$50% higher S.U.N rate than the best previous method.
On the proxy metrics, PLaID++ achieves performance on par with other methods. Our Wyckoff encoding specifically increases compositional validity for both the base and DPO fine-tuned models, highlighting symmetry-based encodings as a natural and intuitive mechanism to increase the validity and stability of generated structures. Though many of these metrics have saturated, we report PLaID++’s performance for comparison and completeness.
We also observe that PLaID++ is significantly more computationally efficient than comparable methods. On a singular NVIDIA 80GB H100 GPU, PLaID++ samples 10,000 crystals in approximately 23 minutes, yielding 27.17 S.U.N. crystals per minute. This generation speed is 5x faster than FlowLLM, which takes 89.6 minutes to generate 10,000 materials on a NVIDIA 80GB A100 GPU, yielding 5.25 S.U.N. crystals per minute (Sriram et al., 2024b). Such high throughput is essential in real-world discovery, as faster sampling dramatically shortens the design–build–test cycle, enabling rapid screening of candidates and speedups of experimental pipelines (Szymanski et al., 2023; Curtarolo et al., 2013).
Conditional Generation   To test the ability of PLaID++ to generalize towards targeted structure synthesis, we measure the S.S.U.N. percentage for four model variants: (i) the 3D coordinate fine-tuned Qwen model (ii) the Wyckoff fine-tuned Qwen model, (iii) DPO conditioned only on space group pairs with no dynamic temperature or novelty reward,(iv) DPO trained jointly on space group, and stable and novel pairs, with dynamic temperature sampling (flagship PLaID++). As shown in Figure 4, applying DPO to space-group pairs increases S.S.U.N. by an average of 22% compared to the base Wyckoff model. Incorporating additional preference data from unconditional generation further boosts performance, resulting in a total increase of 47% over the base model—more than doubling the improvement from space-group conditioning data alone.
Groups with low training data occurrences like ($Amm2$, $Im2,P6_{3}/mmc$) show little improvement or sometimes slight degradation after DPO. This supports the elicitation hypothesis of reinforcement learning (Lambert, 2023), whereby RL amplifies existing valuable behaviors in the base model rather than teaching models new ones. With too few examples, neither the space-group nor the stability objective can generate a meaningful training signal for the model to learn from.
Figure 4: Histogram of S.S.U.N. results across models for space group conditional generation tasks.
Reinforcement Learning from Interatomic Potentials (RLIP)   Across multiple iterations of Direct Preference Optimization (DPO), we observe in Figure 3 that DPO successively improves the number of S.U.N. materials generated by Qwen-7B, demonstrating that reinforcement learning is an effective and robust method to improve generation performance. On unconditional and space group conditioned generation, we improve S.U.N. by $\sim$115% and $\sim$50% respectively compared to fine tuning alone. We believe this method could be scaled further given additional compute.
From our ablation on temperature sampling in Figure 3, we highlight that removing dynamic temperature updates across DPO iterations leads to a loss of diversity, where the S.U.N. rate degrades even as stability continues to increase. This suggests that increasing temperature acts as an entropy regularizer which counteracts model collapse and encourages exploration.
Joint Training Improves Performance   Table 1 and Figure 4 show that our DPO model trained together on the unconditional and space group conditioned generation tasks universally outperforms the DPO models trained individually on each respective task. This validates the efficacy of a universal foundational model that can be leveraged across multiple downstream objectives. The performance gains are particularly pronounced in conditional generation, where only a few hundred samples are available in the training set. This suggests that the additional training signal derived from unconditional generation feedback can be successfully transferred to conditional generation.
Space Group Distribution Analysis   We compare visualizations between the space group distribution of the MP-20 dataset we used to perform SFT on our base models, a sample of 10,000 crystals from Qwen SFT’d on the 3D coordinate representation, and a sample of 10,000 crystals we from our PLaID++ model trained on iterative DPO finetuning with our Wyckoff representation in Figure 5. We note that the flagship model favors generating out-of-train-distribution space groups such as $P2_{1}2_{1}2_{1}$ (19) and $Pnma$ (62), a consequence of our RLIP framework’s novelty incentive.
Space groups $Cm$ (8) and $P1$ (1) are prominent in the base 3D coordinate representation PLaID++ model’s crystals, but not in the flagship model. Space group Cm has a single symmetry operation (a glide plane), while space group $P1$ only has the identity operation. In addition, space groups $P6_{3}/mmc$ (194) and $I4mmm$ (139) are among the five most frequently generated space groups in the flagship model, even though they are not in the top five of the training set. This supports the hypothesis that the Wyckoff representation helps PLaID++ generate systems with complex symmetry requirements, skewing the distribution towards generating crystals with higher symmetries and ensuring a more even distribution across all space groups.
Density Functional Theory Performance Validation   Relaxing 1,000 structures using DFT, we find a stability rate of approximately 19.1% and a S.U.N. rate of 13%. These results are consistent with our findings using the eSEN machine learning interatomic potential (Fu et al., 2025), providing strong empirical support for the quality of our generated materials. Notably, the DFT-based S.U.N. rate is higher than that obtained from the full 10,000-sample evaluation using eSEN. This discrepancy is expected due to the smaller DFT sample size, which yields a higher relative uniqueness rate.
Figure 5: A comparison between the three space group distributions allows us to extract valuable insights on how our PLaID variants affect model behavior.
6 Discussion
In this paper, we introduce a novel Wyckoff-based text encoding for crystal structures and introduce a new reinforcement learning framework, Reinforcement Learning from Interatomic Potentials (RLIP), to guide generation towards chemically stable and useful structures. Our method achieves state-of-the-art performance across stability and S.U.N. metrics. Current random structure search methods achieve less than a 1% success rate (Pickard and Needs, 2011) in identifying stable materials—PLaID++ represents a significant acceleration from traditional approaches to future methods with greater stability and real-world utility.
Limitations & Future Work   While PLaID++ achieves state-of-the-art stability and S.U.N. rates, it has several limitations that point toward promising avenues for future work. We have fine-tuned and evaluated exclusively on the MP-20 dataset ($\sim$45K structures), and scaling to larger, more diverse datasets such as Alexandria ($\sim$2.6M structures) (Schmidt et al., 2024) or LeMaterial ($\sim$6.7M structures) (Siron et al., 2024) may further broaden the model’s coverage of chemical space. Our experiments also focus on unconditional and space-group-conditioned generation, but extending generation to other objectives such as band-gap optimization or ionic conductivity targets would demonstrate wider practical utility. We also only explored Direct Preference Optimization (DPO) as our RL method; investigating alternative algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) or Group Relative Policy Optimization (GRPO) (Shao et al., 2024) might uncover different stability–diversity trade-offs.
Impact Statement   Our approach has the potential to dramatically speed up the design of novel materials for renewable energy, electronics, and carbon-capture applications—paving the way for more efficient, sustainable technologies that benefit society. At the same time, deploying generative models carries risks, including the synthesis of dangerous compounds and uneven access to these tools.